OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form. Answer y - 2x = 4 2x - y = -6 2x + y = 6 2x - y = 4

OpenStudy (anonymous):

y - 2 = 2(x - 3) -- distribute through the parenthesis y - 2 = 2x + 6 -- add 2 to both sides y = 2x + 6 + 2 y = 2x + 8 standard form is Ax + By = C y = 2x + 8 -- subtract 2x from both sides -2x + y = 8 -- multiply by -1 to make x positive 2x - y = -8 <== standard form